OpenStudy (anonymous):

In the system shown below, what are the coordinates of the solution that lies in quadrant IV? 2x^2+y^2=33 x^2+y^2+2y=19

OpenStudy (anonymous):

2x^2 + y^2 = 33 x^2 + y^2 + 2y = 19 2x^2 + y^2 = 33 -2(x^2 + y^2 + 2y = 19) 2x^2 + y^2 = 33 -2x^2 - 2y^2 - 4y = -38 -y^2 - 4y = -5 -y^2 - 4y + 5 = 0 y^2 + 4y - 5 = 0 (y + 5)(y - 1) = 0 y = -5, 1 In quadrant IV, the y is negative. Therefore, 1 is not an answer. y = -5 2x^2 + y^2 = 33 2x^2 + (-5)^2 = 33 2x^2 + 25 = 33 2x^2 = 8 x^2 = 4 x = ±4 In quadrant IV, the x is positive. Therefore, -4 is not an answer. x = 4 Therefore, the solution is Quadrant IV is (4, -5)

OpenStudy (samsan9):

it was (2,-5) -_-